OpenStudy (bekkah323):

You roll 2 dice. What is the probability that the sum of the dice is greater than 6 and 1 die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question.

Directrix (directrix):

Use the attached chart of outcomes of tossing a pair of dice to assist.

Directrix (directrix):

@bekkah323

Directrix (directrix):

The blue circle sums are greater than 6. How many of these blues show one die with two spots?

OpenStudy (bekkah323):

4

Directrix (directrix):

There are 36 possible outcomes when tossing a pair of dice. Of these, 4 outcomes have a sum greater than 6 AND one of the two dice is a "2." Probability is the number of favorable outcomes divided by the number of possible outcomes. For this problem, there are 4 outcomes that we want of the 36 possible outcomes. Probability (sum of 6 and one die = 2) = 4/36 = 1/9.

OpenStudy (bekkah323):

Thank You!

Directrix (directrix):

Glad to help. I'm over on your deck of cards problem now. Join me.